OpenStudy (anonymous):

Write an equation in : vector form, parametric form, and symmetric form given (1,2,3) (-3,2,1)

OpenStudy (anonymous):

you mean line equation?

OpenStudy (anonymous):

yes given the points (1,2,3) (-3,2,1)

OpenStudy (anonymous):

ooh that is an easy one

OpenStudy (anonymous):

r=r_0+t*v all except t are vectors

OpenStudy (anonymous):

for symmetric form solve for t and cancel it out.

OpenStudy (anonymous):

symmetric form

OpenStudy (anonymous):

how'd you attach that ?

OpenStudy (anonymous):

vector form: (x,y,z)=(1,2,3) +t(-4,0,-2) parametric form: (x,y,z) = (1-4t,2,3-2t) Symetric: (x-1)/-4=y-2=(z-3)/-2

OpenStudy (anonymous):

there is an attach file button

OpenStudy (anonymous):

okay thank you both !

OpenStudy (anonymous):

yw